(c+2) +8=4
Can you please help me I'm been home with swine flu for the last 5 days, im in the 8th grade taking algebra please help

Answer by checkley77(12844)

(c+2) +8=4
C+2+8=4
C=4-2-8
C=-6 ANS.
PROOF:
-6+2+8=4
-6+10=4
4=4